# Thermal evolution of dark matter in the early universe from a symplectic glueball model

### v1

@@ Acknowledgments
We thank Nicolao Fornengo, Helena Kolešová, and Zhi-Wei Wang for helpful discussions. The numerical simulations were run on machines of the Consorzio Interuniversitario per il Calcolo Automatico dell’Italia Nord Orientale (CINECA) and of the European Organization for Nuclear Research (CERN). We acknowledge support from the SFT Scientific Initiative of INFN. At the beginning of this project, M.B. was supported by the Italian program for young researchers “Rita Levi Montalcini”. The work of N.F. is supported by the UKRI Science and Technology Facilities Council (STFC) grant ST/X508834/1. The work of M.P. has been partially supported by the Italian PRIN “Progetti di Ricerca di Rilevante Interesse Nazionale – Bando 2022”, prot. 2022TJFCYB, CUP D53D23002970006. M.B. and M.P. have been partially supported by the Italian Research Centre in High-Performance Computing, Big Data and Quantum Computing (ICSC), funded by the European Union – NextGenerationEU. The work of A.S. is supported by the STFC consolidated grant No. ST/X000648/1.
Research Data Access Statement—The data generated for this work and the analysis code can be downloaded from refs. [276] and [277], respectively.
Open Access Statement—For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
@@ Appendix A Generalities about the compact symplectic group
For every $N$, the compact symplectic group $\operatorname{\mathrm{Sp}}(N)$ is a simply connected Lie group (and, hence, is its own universal covering group). Other interesting properties include the fact that the center of this group is always $\left\{\operatorname{\mathbbm{1}}_{2N},-\operatorname{\mathbbm{1}}_{2N}\right\}\cong\operatorname{\mathbb{Z}}_{2}$; the fundamental group and the second homotopy group are trivial, $\pi_{1}(\operatorname{\mathrm{Sp}}(N))=\pi_{2}(\operatorname{\mathrm{Sp}}(N))=\operatorname{\mathbb{Z}}_{1}$, while the third homotopy group is the group of integers: $\pi_{3}(\operatorname{\mathrm{Sp}}(N))=\operatorname{\mathbb{Z}}$ [278, 279].
Writing a generic element of the $\operatorname{\mathrm{Sp}}(N)$ group as $U=\exp(iX)$, with $X$ an Hermitian and traceless $2N\times 2N$ matrix (so that $U\in\operatorname{\mathrm{U}}(2N)$), the condition (A.6) implies the additional requirement $X^{\star}=-\Omega H{\Omega}^{\tiny{T}}$. Equivalently, the Lie algebra $\mathrm{sp}(N)$ of the compact symplectic group can be thought of as the set of quaternion-valued $N\times N$ matrices satisfying $X=-X^{\dagger}$ (where the dagger denotes the transpose of the quaternion-conjugate matrix). The $\mathrm{sp}(N)$ algebra has dimension $N(2N+1)$ and rank $N$; it can be encoded in the Dynkin diagram
@@
For every $\mathrm{sp}(N)$, each irreducible representation $R$ is either real (when the sum of the odd components $a_{i}$ of the Dynkin label of $R$ with an odd index $i$ is even) or pseudo-real (when the sum of the odd-index components of the Dynkin label is odd). The dimension of a generic representation $R$ is given by the Weyl formula
@@
and the roots are (±$(\pm 1,0)$, $(0,\pm 1)$, ±$\pm\left(\frac{1}{2},\frac{1}{2}\right)$, ±$\pm\left(\frac{1}{2},-\frac{1}{2}\right)$; the simple roots are $\alpha^{1}=\left(\frac{1}{2},-\frac{1}{2}\right)$ and $\alpha^{2}=(0,1)$, and their coroots are $\alpha^{1\vee}=(2,-2)$ and $\alpha^{2\vee}=(0,2)$. Correspondingly, the highest-weight vectors of the two fundamental representations, with Dynkin labels (1 0)10(1\,0)( 1 0 ) and (0 1)01(0\,1)( 0 1 ), are $\mu^{1}=\left(\frac{1}{2},0\right)$ and $\mu^{2}=\left(\frac{1}{2},\frac{1}{2}\right)$, respectively. Denoting reflections with respect to the straight line orthogonal to simple root $\alpha^{i}$ and going through the origin as $\sigma_{i}$, and noting that $r_{+}=\sigma_{1}\sigma_{2}$ is a rotation by $\frac{\pi}{2}$ (with $r_{+}^{-1}=r_{-}=\sigma_{2}\sigma_{1}$), it is easy to see that the Weyl group $W$ is the dihedral group $D_{4}$.
@@
which expresses the multiplicity of a weight $\nu$ in an irreducible representation with highest-weight vector $\mu$ in terms of a sum over the elements of the Weyl group, with $\operatorname{sgn}(\sigma)$ denoting the signature of $\sigma$ (which is $1$ when $\sigma$ can be written as the product of an even number of $\sigma_{1}$ and $\sigma_{2}$ factors, while it is $-1$ when $\sigma$ is given by the product of an odd number of $\sigma_{1}$ and $\sigma_{2}$ factors), $\rho$ denoting half of the sum of the positive roots, and $\mathcal{P}(\zeta)$ the Kostant partition function, defined as the number of ways to express $\zeta$ as a linear combination of the positive roots with natural coefficients, it is straightforward to derive the weight multiplicities of any irreducible representation.
The smallest irreducible representations of the $\mathrm{sp}(2)$ algebra are listed in table 3, where $\mathcal{C}_{2}(R)$ denotes the eigenvalue of the quadratic Casimir operator (namely, the sum of the squares of the generators in a generic representation $R$ is equal to $\mathcal{C}_{2}(R)\operatorname{\mathbbm{1}}$) and $\lambda_{R}$ is the Dynkin index (defined as the trace of the square of each generator in the representation $R$—with the products of two different generators being traceless). For a more extensive list of irreducible representations of this algebra (up to dimension $94605$), see ref. [175].
——- Dynkin label dimension $\mathcal{C}_{2}(R)$ $\lambda_{R}$ notes (0 0)00(0\,0)( 0 0 ) $1$ 00 00 trivial, real (1 0)10(1\,0)( 1 0 ) $4$ $\frac{5}{4}$ $\frac{1}{2}$ defining, fundamental, pseudo-real (0 1)01(0\,1)( 0 1 ) $5$ $2$ $1$ fundamental, real (2 0)20(2\,0)( 2 0 ) $10$ $3$ $3$ adjoint, real (0 2)02(0\,2)( 0 2 ) $14$ $5$ $7$ real (1 1)11(1\,1)( 1 1 ) $16$ $\frac{15}{4}$ $6$ pseudo-real (3 0)30(3\,0)( 3 0 ) $20$ $\frac{21}{4}$ $\frac{21}{2}$ pseudo-real (0 3)03(0\,3)( 0 3 ) $30$ $9$ $27$ real (2 1)21(2\,1)( 2 1 ) $35$ $6$ $21$ real (4 0)40(4\,0)( 4 0 ) $35^{\prime}$ $8$ $28$ real (1 2)12(1\,2)( 1 2 ) $40$ $\frac{29}{4}$ $29$ pseudo-real (0 4)04(0\,4)( 0 4 ) $55$ $14$ $77$ real (5 0)50(5\,0)( 5 0 ) $56$ $\frac{45}{4}$ $63$ pseudo-real (3 1)31(3\,1)( 3 1 ) $64$ $\frac{35}{4}$ $56$ pseudo-real (1 3)13(1\,3)( 1 3 ) $80$ $\frac{47}{4}$ $94$ pseudo-real (2 2)22(2\,2)( 2 2 ) $81$ $10$ $81$ real (6 0)60(6\,0)( 6 0 ) $84$ $15$ $126$ real (0 5)05(0\,5)( 0 5 ) $91$ $20$ $182$ real (4 1)41(4\,1)( 4 1 ) $105$ $12$ $126$ real (7 0)70(7\,0)( 7 0 ) $120$ $\frac{77}{4}$ $231$ pseudo-real (3 2)32(3\,2)( 3 2 ) $140$ $\frac{53}{4}$ $\frac{371}{2}$ pseudo-real (1 4)14(1\,4)( 1 4 ) $140^{\prime}$ $\frac{69}{4}$ $\frac{483}{2}$ pseudo-real (0 6)06(0\,6)( 0 6 ) $140^{\prime\prime}$ $27$ $378$ real (2 3)23(2\,3)( 2 3 ) $154$ $15$ $231$ real (5 1)51(5\,1)( 5 1 ) $160$ $\frac{63}{4}$ $252$ pseudo-real ——-
@@
(1 0)(1 0)tensor-product1010(1\,0$\otimes$(1\,0)( 1 0 ) ⊗ ( 1 0 ) =(2 0)(0 1)(0 0),absentdirect-sum200100=(2\,0$\oplus$(0\,1$\oplus$(0\,0),= ( 2 0 ) ⊕ ( 0 1 ) ⊕ ( 0 0 ) , (A.18)
(0 1)(0 1)tensor-product0101(0\,1$\otimes$(0\,1)( 0 1 ) ⊗ ( 0 1 ) =(0 2)(2 0)(0 0),absentdirect-sum022000=(0\,2$\oplus$(2\,0$\oplus$(0\,0),= ( 0 2 ) ⊕ ( 2 0 ) ⊕ ( 0 0 ) , (A.19)
(1 0)(0 1)tensor-product1001(1\,0$\otimes$(0\,1)( 1 0 ) ⊗ ( 0 1 ) =(1 1)(1 0),absentdirect-sum1110=(1\,1$\oplus$(1\,0),= ( 1 1 ) ⊕ ( 1 0 ) , (A.20)
(2 0)(2 0)tensor-product2020(2\,0$\otimes$(2\,0)( 2 0 ) ⊗ ( 2 0 ) =(4 0)(2 1)(0 2)(2 0)(1 0)2(0 0),absentdirect-sum4021022010200=(4\,0$\oplus$(2\,1$\oplus$(0\,2$\oplus$(2\,0$\oplus$(1\,0$\oplus$ 2(0% \,0),= ( 4 0 ) ⊕ ( 2 1 ) ⊕ ( 0 2 ) ⊕ ( 2 0 ) ⊕ ( 1 0 ) ⊕ 2 ( 0 0 ) , (A.21)
@@ Appendix B Details of the simulation updates
there exist different, unitarily inequivalent ways to embed it into an $\operatorname{\mathrm{Sp}}(N)$ matrix, which are determined by the two indices $(i,j)$ specifying the rows and columns where the $\operatorname{\mathrm{SU}}(2)$ elements appears:
@@
    $\mathcal{M}_{k,l}=\left\{\begin{array}[]{ll}\alpha&\mbox{when $k=l=i$ or $k=l=j+N$}\\ \alpha^{\star}&\mbox{when $k=l=j$ or $k=l=i+N$}\\ \beta&\mbox{when $k=i$ and $l=j$}\\ -\beta^{\star}&\mbox{when $k=j$ and $l=i$}\\ \beta^{\star}&\mbox{when $k=i+N$ and $l=j+N$}\\ -\beta&\mbox{when $k=j+N$ and $l=i+N$}\\ 1&\mbox{when $k=l$, with $k$ different from $i$, $j$, $i+N$ and $j+N$}\\ 0&\mbox{otherwise}\end{array}\right.;$ (A.12)
    there exist $\frac{N(N-1)}{2}$ families of matrices of this type, which is the number of ways to choose the strictly ordered pair of indices between $1$ and $N$;
@@
    $\mathcal{N}_{k,l}=\left\{\begin{array}[]{ll}\alpha&\mbox{when $k=l=i$, or when $k=l=j-N$}\\ \alpha^{\star}&\mbox{when $k=l=j$, or when $k=l=i+N$}\\ \beta&\mbox{when $k=i$ and $l=j$, or when $k=j-N$ and $l=i+N$}\\ -\beta^{\star}&\mbox{when $k=j$ and $l=i$, or when $k=i+N$ and $l=j-N$}\\ 1&\mbox{when $k=l$, with $k$ different from $i$, $j-N$, $i+N$ and $j$}\\ 0&\mbox{otherwise}\end{array}\right.;$ (A.19)
    again, there exist $\frac{N(N-1)}{2}$ families of matrices of this type: this results from the product of the number of ways to choose the index $i$ between $1$ and $N$, times the number of ways to choose the index $j$ between $N+1$ and $2N$, with the constraint $j-i\neq N$, divided by two, to account for the fact that the resulting matrix is invariant under the interchange of $i$ with $j-N$;
@@
    $\mathcal{R}_{k,l}=\left\{\begin{array}[]{ll}\alpha&\mbox{when $k=l=i$}\\ \beta&\mbox{when $k=i$ and $l=i+N$}\\ \alpha^{\star}&\mbox{when $k=l=i+N$}\\ -\beta^{\star}&\mbox{when $k=i+N$ and $l=i$}\\ 1&\mbox{when $k=l$, with $k$ different from $i$ and $i+N$}\\ 0&\mbox{otherwise}\end{array}\right.;$ (A.26)

### v2

@@ Acknowledgments
We thank Simone Blasi, Pierluca Carenza, Nicolao Fornengo, Mark Hindmarsh, Helena Kolešová, Kari Rummukainen, Gonzalo Villa, and Zhi-Wei Wang for helpful discussions. The numerical simulations were run on machines of the Consorzio Interuniversitario per il Calcolo Automatico dell’Italia Nord Orientale (CINECA) and of the European Organization for Nuclear Research (CERN). We acknowledge support from the SFT Scientific Initiative of INFN. At the beginning of this project, M.B. was supported by the Italian program for young researchers “Rita Levi Montalcini”. The work of N.F. is supported by the UKRI Science and Technology Facilities Council (STFC) grant ST/X508834/1. The work of M.P. has been partially supported by the Italian PRIN “Progetti di Ricerca di Rilevante Interesse Nazionale – Bando 2022”, prot. 2022TJFCYB, CUP D53D23002970006. M.B. and M.P. have been partially supported by the Italian Research Centre in High-Performance Computing, Big Data and Quantum Computing (ICSC), funded by the European Union – NextGenerationEU. The work of A.S. was supported by the STFC consolidated grant No. ST/X000648/1.
Research Data Access Statement. The data generated for this work and the analysis code can be downloaded from refs. [397] and [398], respectively.
Open Access Statement. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
@@ Appendix A Generalities about the compact symplectic group
For every $N$, the compact symplectic group $\operatorname{\mathrm{Sp}}(N)$ is a simply connected Lie group (and, hence, is its own universal covering group). Locally, the $\operatorname{\mathrm{Sp}}(N)$ group manifold can be written as the product of spheres:
$\operatorname{\mathrm{Sp}}(N)=\prod_{k=1}^{N}S^{4k-1},$ (A.13)
so that, in particular, $\operatorname{\mathrm{Sp}}(1)=S^{3}=\operatorname{\mathrm{SU}}(2)$, whereas the local structure of the $\operatorname{\mathrm{Sp}}(2)$ manifold can be described as the $S^{3}\times S^{7}$ product. Other interesting properties include the fact that the center of the $\operatorname{\mathrm{Sp}}(N)$ group is always $\operatorname{\mathbb{Z}}_{2}$; the fundamental group and the second homotopy group are trivial, $\pi_{1}(\operatorname{\mathrm{Sp}}(N))=\pi_{2}(\operatorname{\mathrm{Sp}}(N))=\operatorname{\mathbb{Z}}_{1}$, while the third homotopy group is the group of integers: $\pi_{3}(\operatorname{\mathrm{Sp}}(N))=\operatorname{\mathbb{Z}}$ [399, 400].
Writing a generic element of the $\operatorname{\mathrm{Sp}}(N)$ group as $U=\exp(iX)$, with $X$ an Hermitian and traceless $2N\times 2N$ matrix (so that $U\in\operatorname{\mathrm{SU}}(2N)$), the condition (A.6) implies the additional requirement $X^{\star}=-\Omega X{\Omega}^{\tiny{T}}$. Equivalently, the Lie algebra $\mathrm{sp}(N)$ of the compact symplectic group can be thought of as the set of quaternion-valued $N\times N$ matrices satisfying $X=-X^{\dagger}$ (where the dagger denotes the transpose of the quaternion-conjugate matrix).
The $\mathrm{sp}(N)$ algebra has dimension $N(2N+1)$ and rank $N$; it can be encoded in the Dynkin diagram
@@
For every $\mathrm{sp}(N)$, each irreducible representation $R$ is either real (when the sum of the components $a_{i}$ of the Dynkin label of $R$ with an odd index $i$ is even) or pseudo-real (when the sum of the odd-index components of the Dynkin label is odd). The dimension of a generic representation $R$ is given by the Weyl formula
@@
and the roots are $(\pm 1,0)$, $(0,\pm 1)$, $\pm\left(\frac{1}{2},\frac{1}{2}\right)$, $\pm\left(\frac{1}{2},-\frac{1}{2}\right)$; the simple roots are $\alpha^{1}=\left(\frac{1}{2},-\frac{1}{2}\right)$ and $\alpha^{2}=(0,1)$, and their coroots are $\alpha^{1\vee}=(2,-2)$ and $\alpha^{2\vee}=(0,2)$. Correspondingly, the highest-weight vectors of the two fundamental representations, with Dynkin labels (1 0)(1\,0) and (0 1)(0\,1), are $\mu^{1}=\left(\frac{1}{2},0\right)$ and $\mu^{2}=\left(\frac{1}{2},\frac{1}{2}\right)$, respectively. Denoting reflections with respect to the straight line orthogonal to the simple root $\alpha^{i}$ and going through the origin as $\sigma_{i}$, and noting that $r_{+}=\sigma_{1}\sigma_{2}$ is a rotation by $\pi/2$ (with $r_{+}^{-1}=r_{-}=\sigma_{2}\sigma_{1}$), it is easy to see that the Weyl group $W$ is the dihedral group $D_{4}$.
@@
which expresses the multiplicity of a weight $\nu$ in an irreducible representation with highest-weight vector $\mu$ in terms of a sum over the elements of the Weyl group, with $\operatorname{sgn}(\sigma)$ denoting the signature of $\sigma$ (which is $1$ when $\sigma$ can be written as the product of an even number of $\sigma_{1}$ and $\sigma_{2}$ factors, while it is 1-1 when $\sigma$ is given by the product of an odd number of $\sigma_{1}$ and $\sigma_{2}$ factors), $\rho$ denoting half of the sum of the positive roots, and $\mathcal{P}(\zeta)$ the Kostant partition function, defined as the number of ways to express $\zeta$ as a linear combination of the positive roots with natural coefficients, it is straightforward to derive the weight multiplicities of any irreducible representation.
The smallest irreducible representations of the $\mathrm{sp}(2)$ algebra are listed in table 5, where $\mathcal{C}_{2}(R)$ denotes the eigenvalue of the quadratic Casimir operator (namely, the sum of the squares of the generators in a generic representation $R$ is equal to $\mathcal{C}_{2}(R)\operatorname{\mathbbm{1}}$) and $\lambda_{R}$ is the Dynkin index (defined as the trace of the square of each generator in the representation RR—with the products of two different generators being traceless). For a more extensive list of irreducible representations of this algebra (up to dimension $94605$), see ref. [193].
——- Dynkin label dimension $\mathcal{C}_{2}(R)$ $\lambda_{R}$ notes (0 0)(0\,0) 11 0 0 trivial, real (1 0)(1\,0) 44 $\frac{5}{4}$ $\frac{1}{2}$ defining, fundamental, pseudo-real (0 1)(0\,1) 55 22 11 fundamental, real (2 0)(2\,0) $10$ 33 33 adjoint, real (0 2)(0\,2) $14$ 55 77 real (1 1)(1\,1) $16$ $\frac{15}{4}$ 66 pseudo-real (3 0)(3\,0) 2020 $\frac{21}{4}$ $\frac{21}{2}$ pseudo-real (0 3)(0\,3) $30$ 99 $27$ real (2 1)(2\,1) $35$ 66 2121 real (4 0)(4\,0) $35^{\prime}$ 88 $28$ real (1 2)(1\,2) $40$ $\frac{29}{4}$ $29$ pseudo-real (0 4)(0\,4) $55$ $14$ $77$ real (5 0)(5\,0) $56$ $\frac{45}{4}$ $63$ pseudo-real (3 1)(3\,1) $64$ $\frac{35}{4}$ $56$ pseudo-real (1 3)(1\,3) $80$ $\frac{47}{4}$ $94$ pseudo-real (2 2)(2\,2) $81$ $10$ $81$ real (6 0)(6\,0) $84$ $15$ $126$ real (0 5)(0\,5) $91$ 2020 $182$ real (4 1)(4\,1) $105$ $12$ $126$ real (7 0)(7\,0) $120$ $\frac{77}{4}$ $231$ pseudo-real (3 2)(3\,2) $140$ $\frac{53}{4}$ $\frac{371}{2}$ pseudo-real (1 4)(1\,4) $140^{\prime}$ $\frac{69}{4}$ $\frac{483}{2}$ pseudo-real (0 6)(0\,6) $140^{\prime\prime}$ $27$ $378$ real (2 3)(2\,3) $154$ $15$ $231$ real (5 1)(5\,1) $160$ $\frac{63}{4}$ $252$ pseudo-real ——-
@@
[…]
$(1\,0)\otimes(0\,1)$ $=(1\,1)\oplus(1\,0),$ (A.21)
[…]
@@ Appendix B Details of the simulation updates
there exist different, unitarily inequivalent ways to embed it into an $\operatorname{\mathrm{Sp}}(N)$ matrix, which are determined by the two indices $(i,j)$ specifying the rows and columns where the $\operatorname{\mathrm{SU}}(2)$ element appears:
@@
[…]
    there exist N(N1)/2N(N-1)/2 families of matrices of this type, which is the number of ways to choose the strictly ordered pair of indices between $1$ and $N$;
@@
[…]
    again, there exist N(N1)/2N(N-1)/2 families of matrices of this type: this results from the product of the number of ways to choose the index $i$ between $1$ and $N$, times the number of ways to choose the index $j$ between $N+1$ and $2N$, with the constraint $j-i\neq N$, divided by two, to account for the fact that the resulting matrix is invariant under the interchange of $i$ with jNj-N;
@@
[…]